Study B X Y 4 5 8 7 8 6 13
Use the following data set and calculate the following: ΣΧ. ΣΥ (Σ)(Σ ) а. b. ΣΧΥ c. Pearson r
Answer #1

@ EX = 4+5+ 7+...+3 Ey = 9+8+8+ ...+ 7 = 71 (EX) (EY) = (41)(71) = 2911] O E XY = 4*9 + 5*8+ ... + 3*7 1 = [325] @ n=91 EX²=4